A river has a current of 5 mph. It takes Al half an hour longer to paddle upstream 1.2 miles than to paddle downstream the same distance. What is Al's rate in still water?
I started out with:
r=rate
With current=r+5
against current=r-5
but this is as far as I could get

Answer by Alan3354(69443)

A river has a current of 5 mph. It takes Al half an hour longer to paddle upstream 1.2 miles than to paddle downstream the same distance. What is Al's rate in still water?
---------
Your statements are correct. Use them in the equations.
d = rt (r = Als rate)
t = 1.2/(r+5)
t+1/2 = 1.2/(r-5)
(1.2/(r+5)) + 1/2 = 1.2/(r-5)
Multiply by 2(r+5)*(r-5)
2.4(r-5) + (r+5)(r-5) = 2.4(r+5)
2.4r-12 + r^2-25 = 2.4r + 12
r^2 - 49 = 0
r = 7 mph
------
1.2 miles at 2 mph = 0.6 hours
1.2 miles at 12 mph = 0.1 hours, 1/2 hour less